OpenStudy (anonymous):

Need help with this simple exponent problem please. 125 exponents: 1/3

OpenStudy (anonymous):

yes please. \[125^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

well the instructions say "write all answers with positive exponents" An it starts with the 125 etc.

OpenStudy (anonymous):

no, it starts with 125. What I don't know how to do is work with fraction exponents.

OpenStudy (anonymous):

That is a positive exponent. 125^(1/3) = \[\sqrt[3]{125}\] Which equals 5.

OpenStudy (anonymous):

thanks! got it. Thanks to both.